# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v2

@@ C.1 Thresholds
To extract thresholds for the codes at various noise-bias strengths, we perform a finite-size data collapse [70, 71] by fitting the data to the curve $A+Bx+Cx^{2}$ where $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for the code-capacity and SDEM3 noise models, respectively. Here, $\nu$ is the critical exponent, $p_{\mathrm{th}}$ is the Pauli threshold and $A,B,C$ are the fit parameters. Figures 12, 13, 14 and 15 show the resulting data collapse and close-to-threshold plots for various bias strengths for the X3Z3, CSS, P6 and XYZ2 Floquet codes, respectively.
C.2 Error suppression factor
Code-capacity SDEM3
$\eta=1$ $\eta=99$ $\eta=1$ $\eta=99$
CSS 1.94 1.39 1.74 1.37
XYZ2 2.27 2.28 1.11 1.15
P6 2.31 2.37 1.12 1.24
X3Z3 2.49 4.26 1.94 3.42
Table 1: Error suppression factors $\Lambda$ for different Floquet codes calculated using two different noise models, code-capacity and SDEM3, with varying levels of noise bias $\eta$. Results are calculated with a physical error rate $p=0.55\%$. The lambda factor is related to the error suppresion rate $\gamma$ shown in Fig. 8
In Table 1, we list the values of the error suppression factor $\Lambda$ which are calculated from the error suppression rate $\gamma$ shown in Fig. 8.
Appendix D Hyperedge errors in honeycomb codes under infinite noise bias
Figure 16: Parts of the detector hypergraphs of (a) P6 and (b) XYZ2 honeycomb codes. Coloured horizontal layers indicate type of measurements performed. Coloured nodes are detectors. Hyperedges and edges between detectors correspond to the syndromes of the single-qubit $Z$ errors shown by the uncoloured circles. Solid (dashed) edges correspond to syndromes due to the $Z$ error on the left (right) qubit. The colour of the (hyper)edge denotes the time step during which the $Z$ error occurs, e.g., a red hyperedge is the syndrome for a $Z$ error occurring immediately after red check measurements. Insets show the syndromes for the two-qubit $ZZ$ error occurring on the two qubits indicated (along a green edge) immediately after green check measurements. This syndrome is edge-like (hyperedge-like) for the P6 (XYZ2) code.
Unlike the CSS and X3Z3 Floquet codes, the single-qubit $X$ and $Z$ Pauli errors in both of the honeycomb code variants can produce hyperedge syndromes [12, 22] if they occur at specific measurement subrounds. These hyperedges degrade the performance of a matching decoder as they must be decomposed into edges, thereby losing some information about the correlations between detectors. We show in Fig. 16 an example of such hyperedges which are formed from single-qubit $Z$ errors. Note that for a single-qubit $Z$ error to produce a hyperedge-like syndrome, it must occur between the measurements of the $XX$ and $YY$ edges of the $Z$-error affected qubit.
First, let us consider the P6 honeycomb code. Since in the P6 code, the edge type is determined by its colour, a single-qubit $Z$ error here generates a hyperedge syndrome (four flipped detectors) if it occurs between red ($XX$) and green ($YY$) check measurements [see the hyperedge highlighted in red in Fig. 16(a)] and an edge-like syndrome (two flipped detectors) if it occur at other measurement subrounds [shown by blue/green edges in Fig. 16(a)]. Notice that the product of the two $Z$ errors occurring simultaneously always results in a graph-like syndrome whenever the error occurs [see inset of Fig. 16(a)]. This is the same for all neighbouring $ZZ$ errors occurring along edges in the lattice.
In the XYZ2 honeycomb code, by contrast, hyperedges can be formed from $Z$ errors occurring at every measurement subround since edges of each colour consist of all check types ($XX$, $YY$ and $ZZ$). The hyperedge-forming $Z$ errors, however, need to occur at specific locations such that they happen between the measurements of the $XX$ and $YY$ edges of the noisy qubits. That is, a $Z$ error with coordinates $(x,y,t)$ could form a hyperedge syndrome for any $t$, given suitably chosen $x$ and $y$. Two such $Z$ errors and their syndromes for different time steps are shown in Fig. 16(b). As can be seen, one $Z$ error (the error on the right) produces a hyperedge if it occurs after red check measurements, while the other (the error on the left) does so only after the green measurement subround. As such, a neighbouring $ZZ$ error occurring at a given time always produce a hyperedge syndrome. For example, if the $ZZ$ error shown (supported by a green edge) in Fig. 16(b) occurs after the green MPP measurements, the resulting syndrome will include four detectors [see inset of Fig. 16(b)].
Appendix E X3Z3 Floquet code with twisted boundary conditions
Figure 17: Evolution of logical operators in the X3Z3 Floquet code with a twisted boundary condition. Number-labelled edges at the top and bottom boundaries indicate the twisted boundary condition. The logical operators are shown in yellow and brown non-trivial cycles.
Figure 18: Thresholds ($p_{\mathrm{th}}$) of the X3Z3 Floquet code with twisted (magenta) and untwisted (blue) periodic boundary conditions as a function of noise bias $\eta$ for two different noise models: (a) code-capacity and (b) SDEM3. We show some of the data points used in obtaining the thresholds in Fig. 19. For better visualisation, we fit all curves with quadratic splines.
Figure 19: Logical failure probability $p_{L}$ of the X3Z3 Floquet code with twisted periodic boundary conditions as a function of $x=(p-p_{\mathrm{th}})L^{1/\nu}$. Data shown are for different lattice lengths $L$ (lattice sizes $L\times 3L/2$). The plots are calculated for two different noise models: code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=999$, and (d,h) $\eta=\infty$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by magenta vertical lines. Results for each lattice length $L$ and physical error rate $p$ are averaged over $10^{6}-10^{7}$ number of shots.
In this appendix, we present simulation results obtained for the X3Z3 code with a twisted boundary condition. The twisted boundary refers to the case where the qubits at the top row of the honeycomb lattice are connected to qubits at the bottom row, shifted by two columns (see Fig. 17). In the figure, we also show two (of the four) logical operators. The yellow logical operator, which consists of only $Z$ operators, has a length $L^{2}$/2 for a lattice with a size $L\times 3L/2$. This is longer than the length of the untwisted code logical operator, which is only $L$ (or $2L$) [cf. Fig. 1(c)]. As a result, the performance of the twisted code near infinite bias is expected to be better than the untwisted code since it will be harder for the pure dephasing errors to form such a long logical operator.
To support this argument, we present the thresholds of such a code in Fig. 18, where we compare them to the thresholds of the untwisted code. Data used to obtain the thresholds are presented in Fig. 19. We find that there is a stronger finite-size effect for the twisted code compared to the untwisted one as can be seen from the finite-size collapse plots. As shown in Fig. 18, the twisted code has better thresholds than the untwisted code only in the regime close to infinite bias. For the small noise bias regime ($\eta\lesssim 10$), the thresholds of the twisted code are roughly similar to those of the untwisted code. However, in the intermediate bias regime, the thresholds of the twisted code are lower than those of the untwisted code.
Besides the thresholds, we also compute the logical failure probability of the twisted X3Z3 Floquet code as a function of the lattice length $L$. As can be seen from Fig. 20, at infinitely biased code-capacity noise the logical failure probability of the untwisted code decays as $p_{L}\propto\mathrm{exp}(-\gamma_{1}L)$, whereas that of the twisted code decays as $p_{L}\propto\mathrm{exp}(-\gamma_{2}L^{2}-\gamma_{1}L)$. However, when the SDEM3 noise model is considered, both the twisted and untwisted code have only an exponential decay. We note that $\gamma_{1}$ and $\gamma_{2}$ become larger for smaller subthreshold physical error rates.
The super-exponential and exponential decay dependences of the twisted X3Z3 code’s logical failure probability with respect to the lattice length $L$ under the code-capacity and SDEM3 noise models, respectively, are due to the fact that the code distances scale quadratically and linearly with the lattice length under these two different noise models. From the “shortest graph-like error” generated via Stim [64], we find that the measurement and Pauli errors in the SDEM3 model can conspire to produce logical error strings with lengths that scale linearly with the lattice length. Using Stim, we also calculate the size of the shortest graph-like error (i.e., the code distance) of the twisted X3Z3 code under the infinitely $Z$-biased noise and find the following relations:
$d_{\mathrm{twist}}$ $=\frac{d_{\mathrm{untwist}}^{2}}{2}=\frac{L^{2}}{2},\quad$ (7a)
$d_{\mathrm{eff,twist}}$ $=2d_{\mathrm{eff,untwist}}-1=L-1,$ (7b)
for the code-capacity and SDEM3 noise models, respectively. Here $d_{\mathrm{twist}}$ and $d_{\mathrm{eff,twist}}$ are the code distances for the twisted X3Z3 code under the code-capacity and SDEM3 noise, respectively. Similarly, $d_{\mathrm{untwist}}$ and $d_{\mathrm{eff,untwist}}$ are the code distances defined for the untwisted code, while $L$ is the length of the honeycomb lattice with size $L\times 3L/2$. We note that for $\eta<\infty$ of both noise models, the size of the shortest graph-like error calculated from Stim is the same for both the untwisted and twisted codes.
In addition to the subthreshold scaling, as can be seen from Fig. 20, for a fixed physical error rate under the infinite-bias code-capacity noise, the logical failure probabilities of the twisted code are much lower than those of the untwisted case while they are roughly similar for the case of SDEM3 noise. The reason is because the infinite-bias code-capacity threshold for the twisted code is higher than that of the twisted code while they are the same for the SDEM3 case, as shown in Fig. 19.
Figure 20: Sub-threshold logical failure probabilities $p_{L}$ vs lattice length $L$ for the untwisted (blue) and twisted (magenta) X3Z3 Floquet codes. Results are calculated for two different noise models: (a) code-capacity and (b) SDEM3 noise at infinite bias $\eta=\infty$. They are also computed with various physical error rates $p$. All curves can be fitted to decay functions $f\propto\mathrm{exp}(-\gamma_{2}L^{2}-\gamma_{1}L)$ or $f\propto\mathrm{exp}(-\gamma_{1}L)$ where $\gamma_{1,2}$ are increasing functions of $(p_{\mathrm{th}}-p)$. Each data point is averaged over $10^{5}-10^{10}$ number of shots.
Appendix F Details of the no-go theorems for 1D decoding of dynamical codes
@@
In this appendix, we provide the proofs for the two theorems in Sec. VI of the main text. We begin by first proving the following lemma which is going to be used in the proof of Theorem 1.
@@
To form detectors at both plaquettes, we require that, for all $M\in\mathcal{M}_{t-1}$, $[M,\prod_{D\in\mathcal{D}_{j}}D]=0$, for $j=1,2$. Otherwise the plaquette operator outcomes would be indeterministic. However, from property (a) of Definition 3 of an HDC (see Sec. VI of the main text), for each $D\in\mathcal{D}_{1}$, there is a measurement in $\mathcal{M}_{t-1}$ that anti-commutes with it. Therefore, $\prod_{D\in\mathcal{D}_{j}}D$ must have even overlap with all edge measurements in $\mathcal{M}_{t-1}$, for $j=1,2$. This condition can only be satisfied for both $j=1,2$ if all $M\in\mathcal{M}_{t-1}$ are also on green edges (if they are supported on red edges, there will be one measurement with odd overlap with each $D\in\mathcal{D}_{1}$, and similarly for $D\in\mathcal{D}_{2}$ if they are supported on blue edges). But now both $\mathcal{M}_{t-1}$ and $\mathcal{M}_{t}$ are made up of measurements on green edges, contradicting property (a) of Definition 3. ∎
We now prove Theorem 1, which we restate here:
@@
A honeycomb dynamical code with duration more than $33k_{f}+k_{i}+$ 3 cannot have a 1D-decodable $Z$-detector graph, where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets.
@@
We take an error occurring in time step $t$ to mean that it occurs immediately after the measurements in $\mathcal{M}_{t}$. Suppose there are $T$ time steps and $n$ qubits in the HDC. Let us consider part of the ZDG that contains vertices corresponding to all detectors formed in time steps $t>k_{i}$, but only those edges corresponding to detectable errors, i.e., those occurring in time steps $k_{i}<t<T-k_{f}$. We will call this subgraph $\overline{\mathrm{ZDG}}$. The number of vertices in this subgraph, corresponding to the number of detectors formed, is at most $n(T-k_{i})/6$, since at each time step no detectors can be formed on neighbouring plaquettes (see Lemma 1), and there are $n/6$ such non-neighbouring plaquettes, which correspond to all plaquettes of a single colour, in the honeycomb lattice without boundary. Meanwhile, $Z$ errors generate $n$ edges between each time step, resulting in a total of $n(T-1-k_{i}-k_{f})$ edges in the $\overline{\mathrm{ZDG}}$.
@@
Let us now create a bipartition of the qubits into sets $C$ and $D$ such that no qubits in $C$ are adjacent to one another, and similarly for qubits in $D$. This is possible since the honeycomb lattice is bipartite. Let us consider the subgraph of $\overline{\mathrm{ZDG}}$, denoted by $\overline{\mathrm{ZDG}}_{C}$, that contains only edges corresponding to $Z$ errors on qubits in $C$. The subgraph $\overline{\mathrm{ZDG}}_{C}$ therefore contains $n(T-1-k_{i}-k_{f})/2$ edges. Suppose two vertices are connected in this graph by more than one edge. In such a scenario, we can choose any two from the collection of edges between these vertices. These two edges correspond to a weight-2 error, i.e, an error with support only on two qubits, in $C$, which is undetectable. This error is undetectable because the combination of the two errors, where each flips the values of the same two detectors, triggers no detector.
@@
Consider one such undetectable weight-2 error, $Z_{i,t}Z_{j,t^{\prime}}$, where $Z_{i,t}$ denotes an error that acts on qubit $i$ at time step $t$ and similarly for $Z_{j,t^{\prime}}$. We will show that for $Z_{i,t}Z_{j,t^{\prime}}$ not to be detectable, the two errors must act on the same qubit ($i=j$) but at adjacent time steps ($t=t^{\prime}\pm 1$). To this end, we begin by ruling out other possibilities. First, we show that if the qubits $i,j\in C$ are different then the error $Z_{i,t}Z_{j,t^{\prime}}$ will anti-commute with some future measurements. This, along with the fact that $Z_{i,t}Z_{j,t}$ is not an operator just measured or about to be measured (since $i$ and $j$ are not connected by an edge), means that the error is detectable by property (c) of Definition 3 of an HDC (see Sec. VI of the main text). We begin by analysing the case where the two errors occur at the same time step ($t=t^{\prime}$). Suppose that this error commutes with all $M\in\mathcal{M}_{t+1}$. Since every measurement in $\mathcal{M}_{t+1}$ must have support only on one qubit in $C$, then the two check measurements in $\mathcal{M}_{t+1}$, one with support on qubit $i$ and the other on qubit $j$, must have Pauli $Z$ operator supports on those respective qubits. Now consider the measurement $M^{\prime}\in\mathcal{M}_{t+2}$ with support on qubit $i$. By property (a) of Definition 3, one measurement from the set $\mathcal{M}_{t+2}$ that has a support on qubit $i$ cannot act as a Pauli $Z$ operator on that qubit, since it has to anti-commute with the previous edge measurement from $\mathcal{M}_{t+1}$ that also has a $Z$ support on the same qubit. Therefore, $M^{\prime}$ will anti-commute with $Z_{i,t}Z_{j,t}$ which, according to property (c) of Definition 3, means that the error $Z_{i,t}Z_{j,t}$ is detectable. Similarly, if the errors occur on different qubits but at different time steps, they must also be detectable. To show this, we can argue using the reasons mentioned above that $Z_{i,t}$ must anti-commute with a future measurement in set $\mathcal{M}_{t+1}$ or $\mathcal{M}_{t+2}$. Since qubits $i$ and $j$ are not connected by an edge in the honeycomb lattice, the same measurement, however, cannot anti-commute with $Z_{j,t^{\prime}}$.
@@
So the only possibility for the weight-2 error to be undetectable is for the two errors to occur on the same qubit, but at different time steps, i.e., $Z_{i,t}Z_{i,t^{\prime}}$ (note that if the two errors on qubit $i$ occur at the same time step then they cancel each other out, which is equivalent to there being no error at all). Suppose, without loss of generality, that $t<t^{\prime}$. For the weight-2 error to be undetectable, the measurements in between them must commute with $Z_{i}$. Since consecutive measurements with support on the same qubit anti-commute, the two errors acting on qubit $i$ must be separated only by one measurement time step where the in-between measurement commutes with the two errors.
@@
The above analysis implies that vertices in $\overline{\mathrm{ZDG}}_{C}$ can be connected by only at most two edges. All such edges would have to correspond to errors on the same qubit and if there were more than two, we could find a pair of edges that correspond to an undetectable error $Z_{i,t}Z_{i,t^{\prime}}$ with $t^{\prime}>t+1$, which we have shown above to be impossible. Therefore, at most half of the edges in $\overline{\mathrm{ZDG}}_{C}$ can be removed without changing which vertices are connected. This means that there is a spanning subgraph of $\overline{\mathrm{ZDG}}_{C}$ with at least $\lceil n(T-1-k_{i}-k_{f})/4\rceil$ edges with no vertices in the subgraph connected by more than one edge.
@@
All things considered, this implies that the full ZDG has a subgraph with $\lceil n(T-1-k_{i}-k_{f})/4\rceil$ edges and a number of vertices $\leq n(T-k_{i})/6$, such that no two vertices are connected by more than one edge. For $T>3k_{f}+k_{i}+3$, there are more edges than vertices in this subgraph, which means that it cannot have all vertices with neighbourhoods of size $\leq 2$. Therefore, some vertices in the full ZDG also must have neighbourhoods with size greater than 2. ∎
We finally provide the proof for Theorem 2:
@@
A GDC with a duration at least $22k_{f}+k_{i}+$ 2, where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets, respectively, cannot have a $Z$-detector graph equivalent to a collection of disjoint repetition codes.

### v3

@@ IV.5 Determining the thresholds
To extract thresholds of the codes, we perform a finite-size collapse [78, 79] of the logical failure probability $p_{L}$ data taken for various physical error rates $p$ and code distances. This is done by fitting the data to the curve $A+Bx+Cx^{2}$ where $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for the code-capacity and SDEM3 noise models, respectively. Here, $\nu$ is the critical exponent, $p_{\mathrm{th}}$ is the Pauli threshold, $A,B,C$ are the fit parameters, $d$ and $d_{\mathrm{eff}}$ are code distances under the code-capacity and SDEM3 noise models, respectively. This data is presented in Supplementary Figs. 5-8 of the Supplementary Information [61].
IV.6 No-go theorems for 1D decoding graphs of dynamical codes
Here, we provide formal statements and proofs for the two no-go theorems presented in the Results section. To this end, we begin by providing several definitions.
Definition 1.
Definition 2.
Finally, we define dynamical codes on the honeycomb lattice in the following way:
Definition 3.
A honeycomb dynamical code (HDC), with “duration” $T$ and constant “initial(final)-time boundary offsets” $t_{i}$ ($t_{f}$), is a finite-depth measurement circuit acting on qubits of the honeycomb lattice (without boundary). The sets of measurements in the circuit, $\mathcal{M}_{t}$ ($t\in\{1,\ldots,T\}$), are composed of two-body Pauli measurements along coloured edges (either $r$, $g$, or $b$, for each time step) of the lattice, performed sequentially on a state stabilised by some group $\mathcal{S}$. The HDC obeys the following properties:
  1. (a)
    Overlapping measurements in consecutive time steps anti-commute and are supported on different edges.
  2. (b)
    Detectors (for time steps $t>t_{i}$) are associated with plaquettes in the lattice (they have support only on qubits around their associated plaquette).
  3. (c)
    If a single-qubit or two-qubit error occurring in time steps $t_{i}<t<T-t_{f}$ anti-commutes with future measurements, it is detectable, unless the two-qubit error is the same as an edge operator just measured or to be measured in the next time step.
  4. (d)
    All single-qubit errors in time steps $t_{i}<t<T-t_{f}$ have syndromes of weight $>1$.
Let us first comment on our definition of an HDC. The chosen properties are very natural. The anti-commutation of consecutive measurement subrounds ensures “local reversibility” of the code, so as to preserve the quantum information and its locality  [80] (although note that we do not require the code to have logical qubits in our definition). The requirement that consecutive sets of measurements are supported on different edges is not very restrictive: if two (anti-commuting) measurements act on the same edge in consecutive time steps, we can replace the second with a Clifford gate and commute that to the end of the circuit. This merely changes the bases of subsequent measurements without changing their (anti-)commutation. Since the Clifford gates are unimportant, we can ignore them and hence we end up with a shorter duration $T$. We restrict our attention to errors occurring in the temporal “bulk” of the code, $t_{i}<t<T-t_{f}$. This avoids complications due to errors occurring close to the initial and final-time boundaries where detectors here may be formed in different ways than the bulk detectors. Specifically, the detectors associated with errors at the initial and final time boundaries are obtained by comparing the edge or plaquette measurement values with the initial state and final read-out of the physical qubits, respectively. The exclusion of these final time boundary detectors is also required here since we do not include final read-out measurements in our definition of an HDC. In summary, we consider errors occurring in all “detection cells” [13] that are completed before time step $T$ and begin at time step $t\geq 1$, where a detection cell consists of all the spacetime points at which a detector can identify errors.
We will show that an HDC cannot even have a 1D-decodable ZDG, which implies that its decoding graph cannot be equivalent to a collection of the much simpler repetition code’s graphs. Note that a repetition code graph is not only 1D-decodable but also has a maximum degree of 2, which means there are no double edges between neighbouring vertices. Here, however, we allow for these double edges in our definition of 1D-decodability, and show that this more general property is also impossible. Such double edges naturally arise in Floquet codes, where they correspond to two-qubit undetectable errors (i.e., edge operators just measured or about to be measured).
@@ IV.6 No-go theorems for 1D decoding graphs of dynamical codes
We begin by first proving the following lemma which is going to be used in the proof of Theorem 1.
@@
To form detectors at both plaquettes, we require that, for all $M\in\mathcal{M}_{t-1}$, $[M,\prod_{D\in\mathcal{D}_{j}}D]=0$, for $j=1,2$. Otherwise the plaquette operator outcomes would be indeterministic. However, from property (a) of Definition 3 of an HDC, for each $D\in\mathcal{D}_{1}$, there is a measurement in $\mathcal{M}_{t-1}$ that anti-commutes with it. Therefore, $\prod_{D\in\mathcal{D}_{j}}D$ must have even overlap with all edge measurements in $\mathcal{M}_{t-1}$, for $j=1,2$. This condition can only be satisfied for both $j=1,2$ if all $M\in\mathcal{M}_{t-1}$ are also on green edges (if they are supported on red edges, there will be one measurement with odd overlap with each $D\in\mathcal{D}_{1}$, and similarly for $D\in\mathcal{D}_{2}$ if they are supported on blue edges). But now both $\mathcal{M}_{t-1}$ and $\mathcal{M}_{t}$ are made up of measurements on green edges, contradicting property (a) of Definition 3. ∎
Using the properties of the HDC as defined above, we now present the first main theorem of our paper:
@@
A honeycomb dynamical code with duration more than $33t_{f}+t_{i}+$ 3 cannot have a 1D-decodable $Z$-detector graph, where $t_{i}$ and $t_{f}$ are the initial and final time-boundary offsets.
Since the proof of the theorem is quite long, here we begin by providing the sketch of the proof. We first show that the $Z$-detector graph of an HDC contains many more edges than vertices. This means the graph must contain many cycles. We will then find a subgraph of the ZDG in which no two vertices are connected by more than one edge, where for such a subgraph, the size of the neighbourhood of each vertex is equal to the number of edges incident to it. In particular, we will show that there exists at least one such subgraph that still possesses many more edges than vertices, meaning it is not possible for all vertices in the subgraph to have neighbourhoods of size $\leq 2$ – if that were the case, then there would be at least as many vertices as edges. This in turn implies that the full ZDG must contain some vertices with neighbourhoods of size $>2$.
The full proof of Theorem 1 is as follows.
@@
We take an error occurring in time step $t$ to mean that it occurs immediately after the measurements in $\mathcal{M}_{t}$. Suppose there are $T$ time steps and $n$ qubits in the HDC. Let us consider part of the ZDG that contains vertices corresponding to all detectors formed in time steps $t>t_{i}$, but only those edges corresponding to detectable errors, i.e., those occurring in time steps $t_{i}<t<T-t_{f}$. We will call this subgraph $\overline{\mathrm{ZDG}}$. The number of vertices in this subgraph, corresponding to the number of detectors formed, is at most $n(T-t_{i})/6$, since at each time step no detectors can be formed on neighbouring plaquettes (see Lemma 1), and there are $n/6$ such non-neighbouring plaquettes, which correspond to all plaquettes of a single colour, in the honeycomb lattice without boundary. Meanwhile, $Z$ errors generate $n$ edges between each time step, resulting in a total of $n(T-1-t_{i}-t_{f})$ edges in the $\overline{\mathrm{ZDG}}$.
@@
Let us now create a bipartition of the qubits into sets $C$ and $D$ such that no qubits in $C$ are adjacent to one another, and similarly for qubits in $D$. This is possible since the honeycomb lattice is bipartite. Let us consider the subgraph of $\overline{\mathrm{ZDG}}$, denoted by $\overline{\mathrm{ZDG}}_{C}$, that contains only edges corresponding to $Z$ errors on qubits in $C$. The subgraph $\overline{\mathrm{ZDG}}_{C}$ therefore contains $n(T-1-t_{i}-t_{f})/2$ edges. Suppose two vertices are connected in this graph by more than one edge. In such a scenario, we can choose any two from the collection of edges between these vertices. These two edges correspond to a weight-2 error, i.e, an error with support only on two qubits, in $C$, which is undetectable. This error is undetectable because the combination of the two errors, where each flips the values of the same two detectors, triggers no detector.
@@
Consider one such undetectable weight-2 error, $Z_{i,t}Z_{j,t^{\prime}}$, where $Z_{i,t}$ denotes an error that acts on qubit $i$ at time step $t$ and similarly for $Z_{j,t^{\prime}}$. We will show that for $Z_{i,t}Z_{j,t^{\prime}}$ not to be detectable, the two errors must act on the same qubit ($i=j$) but at adjacent time steps ($t=t^{\prime}\pm 1$). To this end, we begin by ruling out other possibilities. First, we show that if the qubits $i,j\in C$ are different then the error $Z_{i,t}Z_{j,t^{\prime}}$ will anti-commute with some future measurements. This, along with the fact that $Z_{i,t}Z_{j,t}$ is not an operator just measured or about to be measured (since $i$ and $j$ are not connected by an edge), means that the error is detectable by property (c) of Definition 3 of an HDC. We begin by analysing the case where the two errors occur at the same time step ($t=t^{\prime}$). Suppose that this error commutes with all $M\in\mathcal{M}_{t+1}$. Since every measurement in $\mathcal{M}_{t+1}$ must have support only on one qubit in $C$, then the two check measurements in $\mathcal{M}_{t+1}$, one with support on qubit $i$ and the other on qubit $j$, must have Pauli $Z$ operator supports on those respective qubits. Now consider the measurement $M^{\prime}\in\mathcal{M}_{t+2}$ with support on qubit $i$. By property (a) of Definition 3, one measurement from the set $\mathcal{M}_{t+2}$ that has a support on qubit $i$ cannot act as a Pauli $Z$ operator on that qubit, since it has to anti-commute with the previous edge measurement from $\mathcal{M}_{t+1}$ that also has a $Z$ support on the same qubit. Therefore, $M^{\prime}$ will anti-commute with $Z_{i,t}Z_{j,t}$ which, according to property (c) of Definition 3, means that the error $Z_{i,t}Z_{j,t}$ is detectable. Similarly, if the errors occur on different qubits but at different time steps, they must also be detectable. To show this, we can argue using the reasons mentioned above that $Z_{i,t}$ must anti-commute with a future measurement in set $\mathcal{M}_{t+1}$ or $\mathcal{M}_{t+2}$. Since qubits $i$ and $j$ are not connected by an edge in the honeycomb lattice, the same measurement, however, cannot anti-commute with $Z_{j,t^{\prime}}$.
@@
Using the above reasonings, we can infer that the only possibility for the weight-2 error to be undetectable is for the two errors to occur on the same qubit, but at different time steps, i.e., $Z_{i,t}Z_{i,t^{\prime}}$ (note that if the two errors on qubit $i$ occur at the same time step then they cancel each other out, which is equivalent to there being no error at all). Suppose, without loss of generality, that $t<t^{\prime}$. For the weight-2 error to be undetectable, the measurements in between them must commute with $Z_{i}$. Since consecutive measurements with support on the same qubit anti-commute, the two errors acting on qubit $i$ must be separated only by one measurement time step where the in-between measurement commutes with the two errors.
@@
The above analysis implies that vertices in $\overline{\mathrm{ZDG}}_{C}$ can be connected by only at most two edges. All such edges would have to correspond to errors on the same qubit and if there were more than two, we could find a pair of edges that correspond to an undetectable error $Z_{i,t}Z_{i,t^{\prime}}$ with $t^{\prime}>t+1$, which we have shown above to be impossible. Therefore, at most half of the edges in $\overline{\mathrm{ZDG}}_{C}$ can be removed without changing which vertices are connected. This means that there is a spanning subgraph of $\overline{\mathrm{ZDG}}_{C}$ with at least $\lceil n(T-1-t_{i}-t_{f})/4\rceil$ edges with no vertices in the subgraph connected by more than one edge.
@@
All things considered, this implies that the full ZDG has a subgraph with $\lceil n(T-1-t_{i}-t_{f})/4\rceil$ edges and a number of vertices $\leq n(T-t_{i})/6$, such that no two vertices are connected by more than one edge. For $T>3t_{f}+t_{i}+3$, there are more edges than vertices in this subgraph, which means that it cannot have all vertices with neighbourhoods of size $\leq 2$. Therefore, some vertices in the full ZDG also must have neighbourhoods with size greater than 2. ∎
The above theorem implies that no matter how an HDC is bias-tailored, it likely can never achieve as high a threshold as the repetition code or the bias-tailored stabiliser codes. Indeed, for large $T$ ($T\gg t_{i},t_{f}$), there will be at least approximately $6$ times as many edges as vertices in the ZDG, and there is a subgraph with $3/2$ as many edges as vertices without any pair of vertices sharing more than one edge. This means the graph must have many cycles with length at least three edges.
Definition 4.
A general dynamical code (GDC) is a finite-depth measurement circuit acting on qubits associated with vertices of a graph $G$. It has a duration $T$ and initial (final) time boundary offsets $t_{i}$ ($t_{f}$), as defined above. The sets of measurements $\mathcal{M}_{t}$ are two-body Pauli measurements associated with edges of $G$ (each qubit is in the support of one measurement in each time step), performed sequentially on a state stabilised by some group $\mathcal{S}$. A GDC obeys properties (a) and (d) of an HDC.
@@
A GDC with a duration at least $22t_{f}+t_{i}+$ 2 (i.e., a constant), where $t_{i}$ and $t_{f}$ are the initial and final time-boundary offsets, respectively, cannot have a $Z$-detector graph equivalent to a collection of disjoint repetition codes.